OpenStudy (anonymous):

Evaluate the following expression using the values given. Find 4x − 2y if x = −1 and y = −2

OpenStudy (anonymous):

substitute !

OpenStudy (cwrw238):

(4* -1) - (2*-2)

OpenStudy (anonymous):

\[\Large 4(-1)-2(-2)=-4+4=0\]